piston-cylinder contains gas initially at 3500 kPaa with a volume of 0.03 mº. The gas is compressed during a process where pV 1.25 C to a pressure of 8500 kPaa. The heat transfer from the gas is 2.5 kJ. Determine the change in internal energy, neglecting changes in kinetic and potential energies.


1
Expert's answer
2021-12-10T06:17:01-0500

Initial pressure(p1)=3500x10Pa

Initial Volume (V1)=0.03 m3

Final pressure(p2)=8500 x10Pa

Gas undergoes a polytropic process given by pV1.25=constant

Therefore, for process 1-2 we can write,

"p_1 \times {V_1}^{1.25} = p_2 \times {V_2}^{1.25}"

"{V_2}^{1.25} =\frac { p_1 \times {V_1}^{1.25} }{p_2}"

The work done during poly tropic process is given by

"W=\frac {p_1 \times V_1-p_2 \times V_2}{n-1}"

"W=\frac {(3500 \times (0.03)-(8500)\times (0.01475)}{1.25-1}"

Work done on the gas is W=81.5 kJ

Δ𝑈= −2500+81500=79000𝐽=79𝑘𝐽